Breakfasttime Cereal Company manufactures two breakfast cereals in a joint process. Cost and quantity information is as follows:

  

Joint Cost Cereal Quantity at Split-Off Point Sales Price per Kilogram
$ 93,000 Yummies 12,300 kilograms $ 6.30
Crummies 8,300 kilograms 7.80

Breakfasttime Cereal Company has an opportunity to process its Crummies further into a mulch for ornamental shrubs. The additional processing operation costs $1.80 per kilogram, and the mulch will sell for $12.00 per kilogram.

2.

Suppose the company does process Crummies into the mulch. Use the net-realizable-value method to allocate the joint production cost between the mulch and the Yummies. (Round your calculation of relative proportions to 3 decimal places.)

Allocation of Joint Cost
Yummies
Mulch
Answer #1
Net realisable value = Sales Value - Additional processing cost
NRV Joint costs allocated
Yummies           77,490                   44,443.85
Mulch           84,660                   48,556.15
Total        162,150                   93,000.00

A В C 1 Net realisable value Joint costs allocated =93000* B3/B5 =93000* B4/B5 =SUM(C3:C4) NRV 2 -6.3*12300 3 Yummies (12-1.8